# Utilizing and extending superconducting circuit toolbox to simulate analog quantum gravity

### v1

@@ Appendix A Josephson junction array with gyrators
$\mathcal{H}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L}\left(\phi_{j+1}-\phi_{j}\right)^{2}+M\phi_{j}^{2}$ (43)
@@
$N_{k}$ $=\frac{1}{\sqrt{J}}\sum_{j=1}^{J}e^{ik_{m}j}N_{j}$
$\phi_{k}$ $=\frac{1}{\sqrt{J}}\sum_{j=1}^{J}e^{-ik_{m}j}\phi_{j}$ (44)
$k_{m}$ $=\frac{2\pi m}{J}$
@@
$\mathcal{H}=E_{C}\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\left[N_{m}N_{-m}+i2G\sin\left(\frac{2\pi m}{J}\right)\left\{N_{m},\phi_{m}\right\}+\beta_{m}\phi_{m}\phi_{-m}\right]$ (45)
@@
Now we write the Fourier transformed charge and phase operators in terms of Bosonic annihilation and creation operators
$N_{m}$ $=\frac{i}{\sqrt{2}}\left(a_{-m}-a_{m}^{\dagger}\right)$ (46)
$\phi_{m}$ $=\frac{1}{\sqrt{2}}\left(a_{m}+a_{-m}^{\dagger}\right)$ (47)
that obey the following commutation relations $[a_{m},a^{\dagger}_{m^{\prime}}]=\delta_{mm^{\prime}}$, $[a_{m},a_{m^{\prime}}]=0$. Finally the Hamiltonian becomes
$\mathcal{H}$ $=\frac{E_{C}}{2}\sum_{m=1}^{\frac{J}{2}-1}\left(\begin{array}[]{cccc}a_{m}^{\dagger}&a_{-m}^{\dagger}&a_{m}&a_{-m}\end{array}\right)H_{m}\left(\begin{array}[]{c}a_{m}\\ a_{-m}\\ a_{m}^{\dagger}\\ a_{-m}^{\dagger}\end{array}\right)+\frac{E_{C}}{2}\left(\begin{array}[]{cc}a_{0}^{\dagger}&a_{0}\end{array}\right)h\left(\begin{array}[]{c}a_{0}\\ a_{0}^{\dagger}\end{array}\right)+\frac{E_{C}}{2}\left(\begin{array}[]{cc}a_{\frac{J}{2}}^{\dagger}&a_{\frac{J}{2}}\end{array}\right)h^{\prime}\left(\begin{array}[]{c}a_{\frac{J}{2}}\\ a_{\frac{J}{2}}^{\dagger}\end{array}\right)$ (59)
$\mathcal{H}$ $=\sum_{m=1}^{\frac{J}{2}-1}\mathcal{H}_{1,m}+\mathcal{H}_{2}+\mathcal{H}_{3}$ (60)
@@
$H_{m}$ $=\left(\beta_{m}+1\right)\mathbb{I}_{4}+4G\sin\left(\frac{2\pi m}{J}\right)\left(\begin{array}[]{cc}\sigma_{z}&0\\ 0&\sigma_{z}\end{array}\right)+\left(\beta_{m}-1\right)\left(\begin{array}[]{cc}0&\sigma_{x}\\ \sigma_{x}&0\end{array}\right)$ (65)
$h$ $=\left(\beta_{0}+1\right)\mathbb{I}_{2}+\left(\beta_{0}-1\right)\sigma_{x}$ (66)
$h^{\prime}$ $=\left(\beta_{\frac{J}{2}}+1\right)\mathbb{I}_{2}+\left(\beta_{\frac{J}{2}}-1\right)\sigma_{x},$ (67)
@@
$\mathcal{H}_{1,m}$ $=\frac{E_{C}}{2}\left(\begin{array}[]{cccc}b_{m}^{\dagger}&b_{-m}^{\dagger}&b_{m}&b_{-m}\end{array}\right)\tau_{z}\Lambda_{m}\left(\begin{array}[]{c}b_{m}\\ b_{-m}\\ b_{m}^{\dagger}\\ b_{-m}^{\dagger}\end{array}\right)$ (73)
@@
$=\frac{E_{C}}{2}\left(\begin{array}[]{cccc}b_{m}^{\dagger}&b_{-m}^{\dagger}&b_{m}&b_{-m}\end{array}\right)\tau_{z}\Lambda_{m}\left(\begin{array}[]{c}b_{m}\\ b_{-m}\\ b_{m}^{\dagger}\\ b_{-m}^{\dagger}\end{array}\right)$ (89)
@@
$\mathcal{H}$ $=\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\omega_{m}b_{m}^{\dagger}b_{m}$ (90)
$\omega_{m}$ $=2\sqrt{E_{C}^{2}\beta_{m}}+4E_{C}G\sin\left(\frac{2\pi m}{J}\right)$ (91)
Insert graph here
@@
$\mathcal{H}^{\prime}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L}\left(\phi_{j+1}-\phi_{j-1}\right)^{2}+M\phi_{j}^{2}$ (92)
@@
$\mathcal{H}^{\prime}$ $=\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\omega_{m}b_{m}^{\dagger}b_{m}$ (93)
$\omega_{m}$ $=2\sqrt{E_{C}^{2}\beta_{m}}+4E_{C}G\sin\left(\frac{2\pi m}{J}\right)$ (94)
$\beta_{m}$ $=\frac{4\left(E_{C}G^{2}+E_{L}\right)\sin^{2}\left(\frac{2\pi m}{J}\right)+M}{E_{C}}$
@@ Appendix B Bogoliubov transformation for calculating correlations after quench
$\mathcal{H}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L,j}\left(\phi_{j+1}-\phi_{j}\right)^{2}+M\phi_{j}^{2}$ (95)
@@
$N_{j}$ $=\frac{i}{\sqrt{2}}\left(a_{j}-a_{j}^{\dagger}\right)$ (96)
$\phi_{j}$ $=\frac{1}{\sqrt{2}}\left(a_{j}+a_{j}^{\dagger}\right)$ (97)
@@
$\mathcal{H}=\mathbf{a}^{\dagger}H\mathbf{a}$ (98)
@@
Here,
$Z=\left(\begin{array}[]{cc}\mathbb{I}_{J}&0\\ 0&-\mathbb{I}_{J}\end{array}\right).$
@@
$E_{L,j}=\begin{cases}-E_{L}&j_{0}<j<j_{1}\\ E_{L}&\text{everywhere else}\end{cases}$ (101)
@@
$P^{-1}$ $=ZXP^{T}XZ$ (108)
$\Lambda_{>}$ $=-X\Lambda_{>}X$ (109)
$\Lambda_{>}$ $=M^{-1}\Lambda^{*}_{>}M$ (110)
@@
$=\sum_{j=1}^{J}\left(\lambda_{>,j}c^{\bullet}_{j}c_{j}+\lambda_{>,j}c_{j}c^{\bullet}_{j}\right)$ (112)
@@
$F(t,t^{\prime})\equiv\left<\mathbf{a}(t)\mathbf{a}^{\dagger}(t^{\prime})\right>$ (113)
@@
$F(t,t^{\prime})=\left(\begin{array}[]{cc}A(t,t^{\prime})&B(t,t^{\prime})\\ C(t,t^{\prime})&D(t,t^{\prime})\end{array}\right)$ (116)
@@
$F_{<}(t,t^{\prime})=Q\left<\mathbf{b}(t)\mathbf{b}^{\dagger}(t^{\prime})\right>ZQ^{-1}Z$ (117)
@@
$F_{<}(t,t^{\prime})=Qe^{-i2(t-t^{\prime})\Lambda_{<}}\frac{\mathbb{I}_{2J}+Z}{2}Q^{-1}Z$ (118)
@@
$=Pe^{-i2\Lambda_{>}t}P^{-1}Q\frac{\mathbb{I}_{2J}+Z}{2}Q^{-1}Pe^{i2\Lambda_{>}t^{\prime}}P^{-1}Z$ (119)
@@ Appendix C Klich’s determinant formula
$\hat{A}$ $=\sum_{i,j}\left<i\right|A\left|j\right>d^{\dagger}_{i}d_{j}$
$\hat{B}$ $=\sum_{i,j}\left<i\right|B\left|j\right>d^{\dagger}_{i}d_{j}$
where $A$ and $B$ are first the quantized operators and the states $\left|i\right>$ span the corresponding single particle Hilbert space. Then it can be shown that insert reference here
$\text{Tr}\left(e^{\hat{A}}e^{\hat{B}}\right)=\text{det}\left(1-\xi e^{A}e^{B}\right)^{-\xi}$ (120)
@@
$\hat{A}=\sum_{i,j=1}^{J}a_{i}^{\dagger}A^{(0)}_{ij}a_{j}+\frac{1}{2}a_{i}^{\dagger}A^{(1)}_{ij}a_{j}^{\dagger}+\frac{1}{2}a_{i}A^{(2)}_{ij}a_{j}=\frac{1}{2}\mathcal{A}-\frac{1}{2}\text{tr}A^{(0)}$ (121)
@@
$\mathcal{A}$ $=\mathbf{a}^{\dagger}A\mathbf{a}$
$A$ $=\begin{pmatrix}A^{(0)}&A^{(1)}\\ A^{(2)}&{A^{(0)}}^{T}\end{pmatrix}$
@@
$\text{Diag}\Lambda=\left(\begin{array}[]{cccccc}\lambda_{1}&\cdots&\lambda_{J}&-\lambda_{J}&\cdots&-\lambda_{1}\end{array}\right)$ (125)
@@
$\implies\left|\lambda\right>^{T}ZX\left(ZA+\lambda\right)^{m}$ $=0$ (126)
@@
$\left|\lambda_{i}\right>^{\dagger}\left|\lambda_{j}\right>$ $=\delta_{ij}$ (129)
@@
$=\left(\begin{array}[]{cccccc}c_{1}^{\bullet}&\cdots&c_{J}^{\bullet}&c_{J}&\cdots&c_{1}\end{array}\right)Z\Lambda\left(\begin{array}[]{cccccc}c_{1}&\cdots&c_{J}&c_{J}^{\bullet}&\cdots&c_{1}^{\bullet}\end{array}\right)^{T}$ (132)
@@
$\mathcal{A}=\sum_{i=1}^{J}\lambda_{i}\left(c_{i}^{\bullet}c_{i}+c_{i}c_{i}^{\bullet}\right)+\sum_{i,j=1}^{J}N_{ij}c_{i}^{\bullet}c_{j}^{\bullet}+\sum_{i<j,i=1}^{J}R_{ij}c_{i}^{\bullet}c_{j}$ (133)
@@
$0\right|_{L}c^{\bullet}_{i}$ $=0\text{ }\forall i$ (134)
$c_{i}\left|0\right>_{R}$ $=0\text{ }\forall i$ (135)
@@
$\left|n_{1},n_{2},\cdots,n_{J}\right>$ $=\frac{{c^{\bullet}_{1}}^{n_{1}}{c^{\bullet}_{2}}^{n_{2}}\cdots{c^{\bullet}_{J}}^{n_{J}}}{\sqrt{n_{1}!n_{2}!\cdots n_{J}!}}\left|0\right>_{R}$ (136)
$n_{1},n_{2},\cdots,n_{J}\right|\right.$ =$0\right|_{L}\frac{{c_{1}}^{n_{1}}{c_{2}}^{n_{2}}\cdots{c_{J}}^{n_{J}}}{\sqrt{n_{1}!n_{2}!\cdots n_{J}!}}$ (137)
@@
$n_{1},n_{2},\cdots,n_{J}\right|\right.\left|m_{1},m_{2},\cdots,m_{J}\right>=\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\cdots\delta_{n_{J},m_{J}}$ (138)
@@
^$\hat{\mathcal{I}}=\sum_{n_{1}=0}^{\infty}\cdots\sum_{n_{J}=0}^{\infty}\left|n_{1},n_{2},\cdots,n_{J}\right>\left<\left<n_{1},n_{2},\cdots,n_{J}\right|\right.$ (139)
@@
$=\prod_{i=1}^{J}\frac{e^{\lambda_{i}/2}}{1-e^{\lambda_{i}}}$ (140)
@@
$\text{Tr}e^{\hat{A}}=e^{-\text{tr}A^{(0)}/2}\det(1-UU^{\dagger}Ze^{ZA/2})^{-1}$ (141)
@@
$\left[\text{Tr}e^{\hat{A}}\right]^{2}=\det(Z)e^{-\text{tr}A^{(0)}}\det(1-e^{ZA})^{-1}$ (142)
@@
$\left[\frac{1}{2}\mathcal{A},\frac{1}{2}\mathcal{B}\right]$ $=\frac{1}{2}\mathcal{C}$ (143)
@@
$\mathcal{C}=\mathbf{a}^{\dagger}Z\left[ZA,ZB\right]\mathbf{a}$
@@
$\left[\text{Tr}\left(e^{\hat{A}}e^{\hat{B}}\right)\right]^{2}=\det(Z)e^{-\text{tr}(A^{(0)}+B^{(0)})}\det(1-e^{ZA}e^{ZB})^{-1}$ (144)

### v2

@@ Appendix A Josephson junction array with gyrators
$\mathcal{H}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L}\left(\phi_{j+1}-\phi_{j}\right)^{2}+M\phi_{j}^{2}\ ,$ (43)
@@
$N_{k}$ $=\frac{1}{\sqrt{J}}\sum_{j=1}^{J}e^{ik_{m}j}N_{j}\ ,$
$\phi_{k}$ $=\frac{1}{\sqrt{J}}\sum_{j=1}^{J}e^{-ik_{m}j}\phi_{j}\ ,$ (44)
$k_{m}$ $=\frac{2\pi m}{J}\ ,$
@@
$\mathcal{H}=E_{C}\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\left[N_{m}N_{-m}+i2G\sin\left(\frac{2\pi m}{J}\right)\left\{N_{m},\phi_{m}\right\}+\beta_{m}\phi_{m}\phi_{-m}\right]\ ,$ (45)
@@
Now, we write the Fourier transformed charge and phase operators in terms of bosonic annihilation and creation operators
$N_{m}$ $=\frac{i}{\sqrt{2}}\left(a_{-m}-a_{m}^{\dagger}\right)\ ,$ (46)
$\phi_{m}$ $=\frac{1}{\sqrt{2}}\left(a_{m}+a_{-m}^{\dagger}\right)\ ,$ (47)
that obey the following commutation relations $[a_{m},a^{\dagger}_{m^{\prime}}]=\delta_{mm^{\prime}}$, $[a_{m},a_{m^{\prime}}]=0$. Finally, the Hamiltonian becomes
$\mathcal{H}$ $=\frac{E_{C}}{2}\sum_{m=1}^{\frac{J}{2}-1}\left(\begin{array}[]{cccc}a_{m}^{\dagger}&a_{-m}^{\dagger}&a_{m}&a_{-m}\end{array}\right)H_{m}\left(\begin{array}[]{c}a_{m}\\ a_{-m}\\ a_{m}^{\dagger}\\ a_{-m}^{\dagger}\end{array}\right)+\frac{E_{C}}{2}\left(\begin{array}[]{cc}a_{0}^{\dagger}&a_{0}\end{array}\right)h\left(\begin{array}[]{c}a_{0}\\ a_{0}^{\dagger}\end{array}\right)+\frac{E_{C}}{2}\left(\begin{array}[]{cc}a_{\frac{J}{2}}^{\dagger}&a_{\frac{J}{2}}\end{array}\right)h^{\prime}\left(\begin{array}[]{c}a_{\frac{J}{2}}\\ a_{\frac{J}{2}}^{\dagger}\end{array}\right)\ ,$ (59)
$\mathcal{H}$ $=\sum_{m=1}^{\frac{J}{2}-1}\mathcal{H}_{1,m}+\mathcal{H}_{2}+\mathcal{H}_{3}\ ,$ (60)
@@
$H_{m}$ $=\left(\beta_{m}+1\right)\mathbb{I}_{4}+4G\sin\left(\frac{2\pi m}{J}\right)\left(\begin{array}[]{cc}\sigma_{z}&0\\ 0&\sigma_{z}\end{array}\right)+\left(\beta_{m}-1\right)\left(\begin{array}[]{cc}0&\sigma_{x}\\ \sigma_{x}&0\end{array}\right)\ ,$ (65)
$h$ $=\left(\beta_{0}+1\right)\mathbb{I}_{2}+\left(\beta_{0}-1\right)\sigma_{x}\ ,$ (66)
$h^{\prime}$ $=\left(\beta_{\frac{J}{2}}+1\right)\mathbb{I}_{2}+\left(\beta_{\frac{J}{2}}-1\right)\sigma_{x}.$ (67)
@@
$\mathcal{H}_{1,m}$ $=\frac{E_{C}}{2}\left(\begin{array}[]{cccc}b_{m}^{\dagger}&b_{-m}^{\dagger}&b_{m}&b_{-m}\end{array}\right)\tau_{z}\Lambda_{m}\left(\begin{array}[]{c}b_{m}\\ b_{-m}\\ b_{m}^{\dagger}\\ b_{-m}^{\dagger}\end{array}\right)\ ,$ (73)
@@
$=\frac{E_{C}}{2}\left(\begin{array}[]{cccc}b_{m}^{\dagger}&b_{-m}^{\dagger}&b_{m}&b_{-m}\end{array}\right)\tau_{z}\Lambda_{m}\left(\begin{array}[]{c}b_{m}\\ b_{-m}\\ b_{m}^{\dagger}\\ b_{-m}^{\dagger}\end{array}\right).$ (89)
@@
$\mathcal{H}$ $=\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\omega_{m}b_{m}^{\dagger}b_{m},$ (90)
$\omega_{m}$ $=2\sqrt{E_{C}^{2}\beta_{m}}+4E_{C}G\sin\left(\frac{2\pi m}{J}\right).$ (91)
@@
$\mathcal{H}^{\prime}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L}\left(\phi_{j+1}-\phi_{j-1}\right)^{2}+M\phi_{j}^{2}.$ (92)
@@
$\mathcal{H}^{\prime}$ $=\sum_{m=-\frac{J}{2}+1}^{\frac{J}{2}}\omega_{m}b_{m}^{\dagger}b_{m},$ (93)
$\omega_{m}$ $=2\sqrt{E_{C}^{2}\beta_{m}}+4E_{C}G\sin\left(\frac{2\pi m}{J}\right),$ (94)
$\beta_{m}$ $=\frac{4\left(E_{C}G^{2}+E_{L}\right)\sin^{2}\left(\frac{2\pi m}{J}\right)+M}{E_{C}}.$
@@ Appendix B Bogoliubov transformation for calculating correlations after quench
$\mathcal{H}$ $=\sum_{j=1}^{J}\Big{[}N_{j}+G(\phi_{j+1}-\phi_{j-1})\Big{]}^{2}+E_{L,j}\left(\phi_{j+1}-\phi_{j}\right)^{2}+M\phi_{j}^{2}.$ (95)
@@
$N_{j}$ $=\frac{i}{\sqrt{2}}\left(a_{j}-a_{j}^{\dagger}\right),$ (96)
$\phi_{j}$ $=\frac{1}{\sqrt{2}}\left(a_{j}+a_{j}^{\dagger}\right),$ (97)
@@
$\mathcal{H}=\mathbf{a}^{\dagger}H\mathbf{a},$ (98)
@@
Here
$Z=\left(\begin{array}[]{cc}\mathbb{I}_{J}&0\\ 0&-\mathbb{I}_{J}\end{array}\right),$
@@
$E_{L,j}=\begin{cases}-E_{L}&j_{0}<j<j_{1},\\ E_{L}&\text{everywhere else,}\end{cases}$ (101)
@@
$P^{-1}$ $=ZXP^{T}XZ,$ (108)
$\Lambda_{>}$ $=-X\Lambda_{>}X,$ (109)
$\Lambda_{>}$ $=M^{-1}\Lambda^{*}_{>}M,$ (110)
@@
$=\sum_{j=1}^{J}\left(\lambda_{>,j}c^{\bullet}_{j}c_{j}+\lambda_{>,j}c_{j}c^{\bullet}_{j}\right),$ (112)
@@
$F(t,t^{\prime})\equiv\left<\mathbf{a}(t)\mathbf{a}^{\dagger}(t^{\prime})\right>,$ (113)
@@
$F(t,t^{\prime})=\left(\begin{array}[]{cc}A(t,t^{\prime})&B(t,t^{\prime})\\ C(t,t^{\prime})&D(t,t^{\prime})\end{array}\right),$ (116)
@@
$F_{<}(t,t^{\prime})=Q\left<\mathbf{b}(t)\mathbf{b}^{\dagger}(t^{\prime})\right>ZQ^{-1}Z.$ (117)
@@
$F_{<}(t,t^{\prime})=Qe^{-i2(t-t^{\prime})\Lambda_{<}}\frac{\mathbb{I}_{2J}+Z}{2}Q^{-1}Z.$ (118)
@@
$=Pe^{-i2\Lambda_{>}t}P^{-1}Q\frac{\mathbb{I}_{2J}+Z}{2}Q^{-1}Pe^{i2\Lambda_{>}t^{\prime}}P^{-1}Z.$ (119)
@@ Appendix C Klich’s determinant formula
$\hat{A}$ $=\sum_{i,j}\left<i\right|A\left|j\right>d^{\dagger}_{i}d_{j},$
$\hat{B}$ $=\sum_{i,j}\left<i\right|B\left|j\right>d^{\dagger}_{i}d_{j},$
where $A$ and $B$ are first the quantized operators and the states $\left|i\right>$ span the corresponding single particle Hilbert space. Then it can be shown that [120]
$\text{Tr}\left(e^{\hat{A}}e^{\hat{B}}\right)=\text{det}\left(1-\xi e^{A}e^{B}\right)^{-\xi},$ (120)
@@
$\hat{A}=\sum_{i,j=1}^{J}a_{i}^{\dagger}A^{(0)}_{ij}a_{j}+\frac{1}{2}a_{i}^{\dagger}A^{(1)}_{ij}a_{j}^{\dagger}+\frac{1}{2}a_{i}A^{(2)}_{ij}a_{j}=\frac{1}{2}\mathcal{A}-\frac{1}{2}\text{tr}A^{(0)},$ (121)
@@
$\mathcal{A}$ $=\mathbf{a}^{\dagger}A\mathbf{a},$
$A$ $=\begin{pmatrix}A^{(0)}&A^{(1)}\\ A^{(2)}&{A^{(0)}}^{T}\end{pmatrix}.$
@@
$\text{Diag}\Lambda=\left(\begin{array}[]{cccccc}\lambda_{1}&\cdots&\lambda_{J}&-\lambda_{J}&\cdots&-\lambda_{1}\end{array}\right),$ (125)
@@
$\implies\left|\lambda\right>^{T}ZX\left(ZA+\lambda\right)^{m}$ $=0,$ (126)
@@
$\left|\lambda_{i}\right>^{\dagger}\left|\lambda_{j}\right>$ $=\delta_{ij}.$ (129)
@@
$=\left(\begin{array}[]{cccccc}c_{1}^{\bullet}&\cdots&c_{J}^{\bullet}&c_{J}&\cdots&c_{1}\end{array}\right)Z\Lambda\left(\begin{array}[]{cccccc}c_{1}&\cdots&c_{J}&c_{J}^{\bullet}&\cdots&c_{1}^{\bullet}\end{array}\right)^{T},$ (132)
@@
$\mathcal{A}=\sum_{i=1}^{J}\lambda_{i}\left(c_{i}^{\bullet}c_{i}+c_{i}c_{i}^{\bullet}\right)+\sum_{i,j=1}^{J}N_{ij}c_{i}^{\bullet}c_{j}^{\bullet}+\sum_{i<j,i=1}^{J}R_{ij}c_{i}^{\bullet}c_{j},$ (133)
@@
$0\right|_{L}c^{\bullet}_{i}$ $=0\text{ }\forall i,$ (134)
$c_{i}\left|0\right>_{R}$ $=0\text{ }\forall i.$ (135)
@@
$\left|n_{1},n_{2},\cdots,n_{J}\right>$ $=\frac{{c^{\bullet}_{1}}^{n_{1}}{c^{\bullet}_{2}}^{n_{2}}\cdots{c^{\bullet}_{J}}^{n_{J}}}{\sqrt{n_{1}!n_{2}!\cdots n_{J}!}}\left|0\right>_{R},$ (136)
$n_{1},n_{2},\cdots,n_{J}\right|\right.$ =$0\right|_{L}\frac{{c_{1}}^{n_{1}}{c_{2}}^{n_{2}}\cdots{c_{J}}^{n_{J}}}{\sqrt{n_{1}!n_{2}!\cdots n_{J}!}}.$ (137)
@@
$n_{1},n_{2},\cdots,n_{J}\right|\right.\left|m_{1},m_{2},\cdots,m_{J}\right>=\delta_{n_{1},m_{1}}\delta_{n_{2},m_{2}}\cdots\delta_{n_{J},m_{J}},$ (138)
@@
^$\hat{\mathcal{I}}=\sum_{n_{1}=0}^{\infty}\cdots\sum_{n_{J}=0}^{\infty}\left|n_{1},n_{2},\cdots,n_{J}\right>\left<\left<n_{1},n_{2},\cdots,n_{J}\right|\right..$ (139)
@@
$=\prod_{i=1}^{J}\frac{e^{\lambda_{i}/2}}{1-e^{\lambda_{i}}}.$ (140)
@@
$\text{Tr}e^{\hat{A}}=e^{-\text{tr}A^{(0)}/2}\det(1-UU^{\dagger}Ze^{ZA/2})^{-1},$ (141)
@@
$\left[\text{Tr}e^{\hat{A}}\right]^{2}=\det(Z)e^{-\text{tr}A^{(0)}}\det(1-e^{ZA})^{-1}.$ (142)
@@
$\left[\frac{1}{2}\mathcal{A},\frac{1}{2}\mathcal{B}\right]$ $=\frac{1}{2}\mathcal{C},$ (143)
@@
$\mathcal{C}=\mathbf{a}^{\dagger}Z\left[ZA,ZB\right]\mathbf{a}.$
@@
$\left[\text{Tr}\left(e^{\hat{A}}e^{\hat{B}}\right)\right]^{2}=\det(Z)e^{-\text{tr}(A^{(0)}+B^{(0)})}\det(1-e^{ZA}e^{ZB})^{-1}.$ (144)